OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x - 2| + 3 = 7 Step 2: |x - 2| = 7 - 3 Step 3: |x - 2| = 4 Which of the following is a correct next step to solve the equation? h of the following is a correct next step to solve the equation? A)x + 2 = -4 B)-x - 2 = 4 C)x + 2 = 4 D)x - 2 = -4

OpenStudy (anonymous):

http://www.purplemath.com/modules/solveabs.htm There are two possible answers. The next step would require removing the absolute value bars and creating two equations. Only one of those is a choice here.

OpenStudy (anonymous):

Solve | x + 2 | = 7 To clear the absolute-value bars, I must split the equation into its two possible two cases, one case for each sign: (x + 2) = 7 or –(x + 2) = 7 x + 2 = 7 or –x – 2 = 7 x = 5 or –9 = x